# Predictively Oriented Posteriors

### v1

@@ 2.1 Exact and Approximate Representations of Predictive Scores
                                                                                                    $\Delta_{\log}(Q,x)$ ${}={}$ $\int\frac{1}{2g_{x}(p_{\theta_{1}},Q)^{2}}\left\{p_{\theta_{1}}(x)-\int p_{\theta_{2}}(x)\mathsf{d}Q(\theta_{2})\right\}^{2}\mathsf{d}Q(\theta_{1}),$                                                                                                      (7)
which is clearly intractable due to the dependence of $g_{x}(p_{\theta_{1}},Q)$ on $\int p_{\theta}(x)\mathsf{d}Q(\theta)$. To approximate $\Delta_{\log}$ in a tractable manner, Masegosa (2020) assumed that $\sup_{\theta\in\Theta,x\in\mathcal{X}}p_{\theta}(x)\leq u$ for some $u>0$, so that $\Delta_{\log}(Q,x)\geq\int\frac{1}{2\sup_{\theta^{\prime}\in\Theta}p_{\theta^{\prime}}(x)}\left\{p_{\theta_{1}}(x)-p_{\theta_{2}}(x)\right\}^{2}\mathsf{d}Q^{2}(\theta_{1:2})$, which leads to a tractable approximation, $\widehat{S}_{\operatorname{DI}}$, that upper-bounds the predictive logarithmic score via
                                                                                                    $S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)\leq\widehat{S}_{\operatorname{DI}}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}:={}$ $\int L_{\operatorname{DI}}(\theta_{1:2},x)\mathsf{d}Q^{2}(\theta_{1:2}),$                                                                                                      (8)
                                                                                                    $L_{\operatorname{DI}}(\theta_{1:2},x)$ ${}:={}$ $-\log p_{\theta_{1}}(x)-\frac{\left\{p_{\theta_{1}}(x)-p_{\theta_{2}}(x)\right\}^{2}}{2\sup_{\theta\in\Theta}p_{\theta}(x)}.$                                                                                                      (9)
Since the term $\Delta_{\log}$ can be viewed as a variance or diversity term, Masegosa (2020) referred to this approximation strategy as diversity inducing. We will pay homage to this work by referring to all tractable predictive scoring rules constructed from the representation in (2.1)—whether exact or approximate—as diversity inducing (DI).
While the approximations suggested by Masegosa (2020) in the log score setting allowed for the derivation of a second-order PAC-Bayes bound, they are not practically appealing: $\sup_{x\in\mathcal{X},\theta\in\Theta}p_{\theta}(x)\leq u$ means that the bound only applies to models whose density or mass function is bounded, which rules out even simple models such as Gaussians. To address this deficiency, Morningstar et al. (2022) proposed a second type of tractable approximation based on multi sample (MS) bounds in variational inference. While originally proposed exclusively for the logarithmic predictive scoring rule, the approach clearly generalises to any convex score. In particular, for any $k\in\mathbb{N}$, we can define the tractable upper bound approximation $\widehat{S}_{\operatorname{MS}}$ as
                                                                                                    $S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)\leq\widehat{S}_{\operatorname{MS}}^{k}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}:={}$ $\int L_{\operatorname{MS}}^{k}(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),$                                                                                                      (10)
                                                                                                    $L_{\operatorname{MS}}^{k}(\theta_{1:k},x)$ ${}:={}$ $S\left(\frac{1}{k}\sum_{j=1}^{k}P_{\theta_{j}},x\right).$                                                                                                      (11)
For larger choices of $k$, this bound tightens, and the approximation error vanishes as $k\to\infty$. For the special case of the logarithmic score, Morningstar et al. (2022) attempted to translate this into a PAC-Bayesian generalisation bound for a PrO posterior. However, the resulting bound was vacuous, and failed to discern whether the resulting posteriors conferred any advantages over Bayes posteriors. We rectify this omission, and deliver such a result: under mild regularity conditions and for arbitrary scoring rules, PrO posteriors based on $\widehat{S}^{k}_{\operatorname{MS}}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ strictly improve predictive performance relative to Bayes posteriors and Gibbs posteriors.
To end this section, we will summarise and generalise our observations on previous approaches through a mild assumption that restricts how the exact or approximate predictive scores giving rise to PrO posteriors will be computed. Importantly, this assumption will ensure that we will not have to differentiate explicitly between the case of exact or approximate predictive scores. Instead, we only assume that the (exact or approximate) predictive score can be computed as an average of form
                                                                                                    $\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}:={}$ $\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),$                                                                                                      (12)
where $L$ is subject to further requirements that generalise both the DI-based approach of Masegosa (2020) and the MS-based approach of Morningstar et al. (2022).
Assumption 2.
$\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)=\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k})$ is the approximate or exact predictive scoring rule, where $L:\Theta^{k}\times\mathcal{X}\to\mathbb{R}$ satisfies conditions $\operatorname{(DI)}$ or $\operatorname{(MS)}$ below.
  • $\operatorname{(DI)}$
    $L(\theta_{1:k},x)=S(P_{\theta_{1}},x)-\delta(\theta_{1:k},x)$, for some $\delta:\Theta^{k}\times\mathcal{X}\to\mathbb{R}_{+}$ and $\Delta(Q,x)\geq\int\delta(\theta_{1:k},x)\mathsf{d}Q(\theta)$, for all $x\in\mathcal{X}$ and $\theta_{1:k}\in\Theta^{k}$. Further, for any $\varphi\in\Theta$, if $\theta_{1:k}=(\varphi,\varphi,\dots\varphi)$, then $\delta(\theta_{1:k},x)=0$, for all $x\in\mathcal{X}$.
  • $\operatorname{(MS)}$
    $L(\theta_{1:k},x)=S(\frac{1}{k}\sum_{j=1}^{k}P_{\theta_{j}},x)$.
2.2 A Key Lemma for PrO Posteriors
Assumption 2 allows us to formally define the PrO posterior irrespective of the sub-case considered, which can now be conveniently represented in the form:
$Q_{n}=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\lambda_{n}\int L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\mathrm{KL}(Q\|\Pi)\right\},$ (13)
where $\lambda_{n}>0$ is a positive scalar,
                                                                                                    $L_{n}(\theta_{1:k})$ ${}={}$ $\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i})$                                                                                                      (14)
denotes the in-sample loss corresponding to the (exact or approximate) score. This representation is the basis for a key Lemma that we build on throughout the remainder, and which is inspired by arguments prevalent for the study of Gibbs posteriors.
Crucially, the theoretical analysis of $Q_{n}$ will deviate sharply from standard approaches due to the appearance of product measures, the generally non-standard loss $L_{n}(\theta_{1:k})$, and the focus on the predictive $P_{Q}:=\int P_{\theta}\;\mathsf{d}Q(\theta)$. Despite these differences, we impose Hoeffding-type moment conditions on the function $L_{n}(\theta_{1:k})$ that are very similar to those arising in the study of Gibbs posteriors.
Assumption 3.
Define the expectation $\theta_{1:k}\mapsto\mathcal{L}(\theta_{1:k}):=\mathbb{E}[L(\theta_{1:k},X)]$. For any $\lambda\in\mathbb{R}$ and some $C_{k}>0$, $\mathop{\text{\large$\int$}}\nolimits\mathbb{E}\left[e^{\lambda\{\mathcal{L}(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\right]\mathsf{d}\Pi^{k}(\theta_{1:k})\leq e^{\lambda^{2}C_{k}^{2}/n}.$
Usually, the above condition is unproblematic: it automatically holds for scoring rules based on the squared MMD with uniformly bounded kernels. For the log score, the situation is more delicate. The log score must always be approximated, so Assumption 3 must be checked for a given approximation strategy. While this is a well-known drawback of imposing Hoeffding-type assumptions in the context of logarithmic scores (see e.g. Germain et al., 2016), it is generally hard to avoid them for this kind of analysis. Indeed, Assumption 3 was also crucial for the developments of both Masegosa (2020) and Morningstar et al. (2022).
With our three main assumptions in place, we can now state our first generic guarantee for PrO posteriors. Recalling that $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$, and $\mathcal{S}(P,P_{0})=\mathbb{E}[S(P,X)]$ with $\mathcal{D}_{S}(P_{Q},P_{0})=\mathcal{S}(P_{Q},P_{0})-\mathcal{S}(P_{0},P_{0})$, we obtain a PAC-Bayes type inequality.
Lemma 1.
Under Assumptions 1-3, the following two statements hold:
                                                                                                    $\mathbb{E}\left[\mathcal{S}(P_{Q_{n}},P_{0})\right]$ ${}\leq{}$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\mathrm{KL}({Q^{k}\|\Pi^{k}})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}{k}C_{k}^{2}}{n};$                                                                                                      (15)
if the scoring rule is exact so that $\widehat{S}=S$, i.e., (6) holds, then
                                                                                                    $\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(P_{Q},P_{0}\right)+\frac{\mathrm{KL}({Q^{k}\|\Pi^{k}})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}{k}C_{k}^{2}}{n}.$                                                                                                      (16)
Lemma 1 delivers a meaningful preliminary bound on the predictive performance of PrO posteriors. Further, this result applies to PrO posterior based on any predictive score—whether approximate or exact—so long as it satisfies Assumptions 1-3. In the specific case of logarithmic score, it also recovers an expectation version of the bounds in both Masegosa (2020) and Morningstar et al. (2022). The implications are immediate. Lemma 1 provides the first known theoretical guarantee for the MMD-based PrO posterior proposed in Shen et al. (2025), and lends theoretical rigour to justify variational approaches seeking to approximate $Q_{n}$ (e.g. Sheth and Khardon, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a; Lai and Yao, 2024).
3 Properties of PrO Posteriors
While Lemma 1 superficially resembles a PAC-Bayes bound used in the study of Gibbs posteriors, this in no way implies that PrO posteriors behave similarly to Gibbs posteriors. In fact, we will show that unlike Gibbs posteriors, PrO posteriors adapt to the level of discord between an assumed model and the observed data.
3.1 Model Misspecification Settings
Before investigating performance and adaptivity of PrO posteriors in more detail, we posit a taxonomy of model misspecification regimes under which we will compare the behaviour of the Gibbs and PrO posteriors.
@@ 3.1 Model Misspecification Settings
To aid understanding, Figure 3 provides an illustration of these different forms of model misspecification. What we call non-trivial misspecification is a generalisation of what Grünwald and van Ommen (2017) termed bad misspecification. It rules out forms of misspecification where there exists a singular model parametrisation that cannot be outperformed by any convex combination of models. Convex recovery is a special case of this, and implies that the extended model class $\{P_{Q}:Q\in\mathcal{P}(\Theta)\}$ obtained via convex combinations of $\mathcal{M}_{\Theta}$ contains $P_{0}$ and is therefore well-specified.
@@
Figure 3: Illustration of key ideas introduced in Definition 2. Panel (a) depicts trivial model misspecification: relative to the chosen score $\mathcal{S}$, a singular element of $\mathcal{M}_{\Theta}$ provides the best fit for $P_{0}$, and there is no benefit to averaging. In contrast, Panel (b) illustrates the case of non-trivial model misspecification: constructing ${Q^{\star}}=\operatornamewithlimits{argmin}_{Q}\mathcal{S}(P_{Q},P_{0})$ as a convex combination of $P_{\theta_{1}}$ and $P_{\theta_{2}}$ yields a better predictive for $P_{0}$ than any one point in $\mathcal{M}_{\Theta}$. Panel (c) provides pictorial intuition for convex recoverability: it is the special case where $P_{Q^{\star}}$ and $P_{0}$ coincide.

### v2

@@ 1.1 Key Ideas
                                                                                                    $\mathsf{d}Q_{n}(\theta)$ ${}\propto{}$ $\mathsf{d}\Pi(\theta)\cdot\exp\left\{-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\left.\frac{\delta S(P_{Q},x_{i})}{\delta Q}\right|_{Q=Q_{n}}(\theta)\right\},$                                                                                                      (5)
where $\frac{\delta S(P_{Q},x)}{\delta Q}\big|_{Q=Q_{n}}$ denotes the functional derivative of $S(P_{Q},x)$ with respect to $Q$ evaluated at $Q={Q_{n}}$ (see also Appendix B.1). While generally of limited use for gaining intuition about its behaviour, this representation illustrates that the PrO posterior is generally not a coherent belief update in the sense of Bissiri et al. (2016).
1.2 Related Work
Approaches and ideas conceptually related to PrO posteriors have appeared in the machine learning literature under a variety of names and motivations. One prominent line of work is rooted in PAC-Bayesian analyses that introduce diversity into an ensemble such as weighted majority voting, whereby a set of classifiers are assigned weights, and the final prediction is the label supported by the largest total weight across the ensemble (Lacasse et al., 2006; Laviolette et al., 2011; Germain et al., 2015; Laviolette et al., 2017; Lorenzen et al., 2019; Masegosa et al., 2020; Viallard et al., 2021; Wu et al., 2021; Wu and Seldin, 2022; Vignault et al., 2023; Theisen et al., 2024). Related PAC-Bayesian perspectives extending beyond majority voting have also been explored by Audibert (2004), Ortega et al. (2022), Hauptvogel and Igel (2024) and Pitas et al. (2025).
More directly related is the recent work of Shen et al. (2025), which proposes the special case of (4) with the MMD in the context of misspecified deterministic simulation models as a heuristic approach “for arresting the collapse of predictive uncertainty” to avoid that “posterior predictions become deterministic.” This is very different from what we study in the remainder: Shen et al. (2025) make no attempt at advocating for the general formulation in (4) as a new general-purpose methodology for parameter and predictive inference, and do not provide any rigorous investigation into its behaviour.
Several alternative heuristic variational approaches targeting objectives similar to (4) have also been suggested elsewhere (see Wang and Liu, 2019; Sheth and Khardon, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a; Lai and Yao, 2024). Amongst these, the most relevant for our remaining developments are two papers studying PAC-Bayesian guarantees for the special case of an approximated logarithmic scoring rule termed $\operatorname{PAC}_{T}^{2}$-variational posterior (Masegosa, 2020) and $\operatorname{PAC}^{m}$-Bayes (Morningstar et al., 2022). Despite superficial similarities, our development differs significantly from these precursors in both form and function. Most critically, this vein of prior work does not define or study the PrO posterior. Instead, the focus is on deriving computationally efficient methods for Bayesian deep learning via variational approximation algorithms informed by objectives similar to (4), as well as the derivation of PAC-Bayes bounds to justify these algorithms.
In summary, previous contributions studying objectives and objects with similarities to (4) are limited in scope, motivated heuristically, and do not explore the promise of PrO posteriors as a new and universally applicable inferential methodology. In the remainder, we rectify this: we provide practically meaningful guarantees for PrO posteriors and study the nature of the uncertainty they communicate. Based on our findings, we argue that PrO posteriors provide a new form of uncertainty quantification, and are uniquely suited to address the predictively oriented inferences that scientists are increasingly demanding from the statistical sciences.
2 A Motivating Example
Throughout, robustness to model misspecification is an important motivation in our development of PrO posteriors. Robustness is also one of the key drivers behind many Gibbs posteriors, usually to guard against Huber contamination (Altamirano et al., 2023a, b; Matsubara et al., 2022, 2023; Laplante et al., 2025; Duran-Martin et al., 2024; Marusic et al., 2025) . Whenever the nature of misspecification is more complicated than can be described through an outlier model of the Huber type, these approaches fail to provide robustness.
To contrast this behaviour with that of PrO posteriors, we consider a linear regression model which for $x_{i}\in\mathbb{R}$, $z_{i}\in\mathbb{R}^{2}$ and $i=1,\dots n$ is given by
                                                                                                    $x_{i}$ ${}={}$ $z_{i}^{\top}\theta+\varepsilon_{i}.$                                                                                                      (6)
Here, we make two assumptions: the parameter of interest $\theta\in\mathbb{R}^{2}$ is fixed but unknown, and $\varepsilon_{i}\overset{\text{iid}}{\sim}\mathcal{N}(0,\sigma^{2})$ for a fixed and known noise level $\sigma^{2}$. Relative to this model for the data, we explore three different forms of misspecification:
  1. (T)
    Heavy-tailed observation noise $\varepsilon_{i}\overset{\text{iid}}{\sim}t_{3}(0,\sigma^{2})$;
  2. (CR)
    Random coefficients $\theta_{2}=2\cdot(-1)^{\xi_{i}}$, with $\xi_{i}\overset{\text{iid}}{\sim}\operatorname{Ber}(1/2)$ sampled from a Bernoulli;
  3. (NT)
    The combined effect of $\varepsilon_{i}\overset{\text{iid}}{\sim}t_{3}(0,\sigma^{2})$ and $\theta_{2}=2\cdot(-1)^{\xi_{i}}$ with $\xi_{i}\overset{\text{iid}}{\sim}\operatorname{Ber}(1/2)$.
Figure 1: The Gibbs and PrO posterior distributions under different forms of model misspecification. Coefficients $\theta_{1},\theta_{2}$ are shown as black crosses, with $\theta_{1}=\theta_{2}$ under trivial misspecification.
The first setting constitutes a form of trivial (T) misspecification: the inflated noise leads to overconfident inferences, but does not compromise the meaning of the model parameter $\theta$. In line with this, we would expect both Gibbs and PrO posteriors to perform reasonably well at identifying the general location of this parameter. In contrast, the second setting fundamentally compromises the interpretation of inferences on $\theta$. However, it allows convex recovery (CR): a 2-component mixture over the misspecified linear regression model recovers the actual data-generating process. While Gibbs posteriors will be adversely affected by this setting, PrO posteriors will be able to recover the true data-generating process exactly. The last setting merges both of the previous complications, and constitutes a form of non-trivial (NT) misspecification: not even convex combinations of the original model class will recover the true data-generating process exactly. Still, we expect that the PrO posterior will perform well, since a convex combination of the original model will get us very close to the true data-generating mechanism.
Figure 2: Mean and standard error of Gibbs posterior’s predictive loss relative to PrO posteriors evaluated using negative log likelihood (NLL), squared MMD ($\operatorname{MMD}^{2}$), and Continuous Ranked Probability Score (CRPS) on a test set for different forms of misspecification.
2.1 Numerical Results
Our numerical results confirm these intuitions, and are summarised in Figures 1, where we compare posterior inferences for Gibbs and PrO posteriors based on the MMD in all three settings, with implementation details and further information deferred to Appendix D.4. Unlike the Gibbs posterior, the PrO posterior performs uncertainty quantification for both convex recovery and the non-trivially misspecified setting by accounting for the distortions of the data-generating mechanism relative to the posited simple linear regression. Even in the trivially misspecified setting, the PrO posterior’s parameter uncertainty seems a more appropriate reflection of the true underlying uncertainty. The reason is that the PrO posterior is driven by its implied predictive distribution, and therefore is more sensitive to the inflated noise in the trivially misspecified setting. Figure 2 translates the impact of parameter uncertainty into predictive performance, and shows an even clearer advantage of the resulting PrO posterior predictive distributions.
2.2 A Taxonomy of Model Misspecification
Based on these observations, it will prove useful to rigorously define this so far intuitively motivated taxonomy of model misspecification. As we shall see later, doing so will prove critical for comparing the behaviours of Gibbs and PrO posteriors.
@@ 2.2 A Taxonomy of Model Misspecification
To aid understanding, Figure 3 provides an illustration of these different forms of model misspecification. What we call non-trivial misspecification is a generalisation of what Grünwald and van Ommen (2017) termed “bad misspecification”. Convex recovery is a special case of this, and implies that the extended model class $\{P_{Q}:Q\in\mathcal{P}(\Theta)\}$ obtained via convex combinations of $\mathcal{M}_{\Theta}$ contains $P_{0}$.
@@
Figure 3: Illustration of key ideas introduced in Definition 2. Panel (a) depicts trivial model misspecification: relative to the chosen score $\mathcal{S}$, a singular element of $\mathcal{M}_{\Theta}$ provides the best fit for $P_{0}$, and there is no benefit to averaging. In contrast, Panel (b) illustrates the case of non-trivial model misspecification: constructing ${Q^{\star}}:=\operatornamewithlimits{argmin}_{Q}\mathcal{S}(P_{Q},P_{0})$ as a convex combination of $P_{\theta_{1}}$ and $P_{\theta_{2}}$ yields a better predictive for $P_{0}$ than any one point in $\mathcal{M}_{\Theta}$. Panel (c) provides pictorial intuition for convex recoverability: it is the special case where $P_{Q^{\star}}$ and $P_{0}$ coincide.
3 Formal Guarantees
The motivating example in the previous section shows empirically that PrO posteriors adapt to model misspecification: when the model is well-specified, they behave like Gibbs or Bayes posteriors, but deliver fundamentally different behaviours when the model is misspecified. In this section, we clarify this adaptivity and formally show that PrO posteriors are predictively superior to Bayes and Gibbs posteriors. Throughout, we will often use the shorthand predictive scoring rule to refer to a scoring rule $S(P_{Q},x)$ applied directly to score a predictive distribution $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$. Additionally, for two sequences $a_{n},b_{n}$ we write $a_{n}\lesssim b_{n}$ to mean that $a_{n}\leq Cb_{n}$ for all $n$ large enough and some universal constant $C>0$. Further, we say that $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
3.1 Regularity conditions
Throughout, we impose a minimal convexity assumption satisfied by virtually all scoring rules of interest—including kernel scoring rules and the logarithmic score.
Assumption 1.
In addition to convexity of the scoring rule, we require regularity conditions on their expectations. While less restrictive conditions can be obtained, the following regularity condition is easy to state, and suffices for our theoretical analysis.
Assumption 2.
Let $r_{n}$ be a positive sequence such that $r_{n}\rightarrow 0$ as $n\rightarrow\infty$ and $\mathcal{P}_{\Pi}(\Theta):=\{Q\in\mathcal{P}(\Theta):Q\ll\Pi\}$, it holds that $\mathbb{E}\left[\sup_{Q\in\mathcal{P}_{\Pi}(\Theta)}\left|\frac{1}{n}\sum_{i=1}^{n}S(P_{Q},X_{i})-\mathcal{S}(P_{Q},P_{0})\right|\right]\lesssim r_{n}$.
As the speed at which the sequence $r_{n}$ decays depends on the scoring rule, we state it in full generality. For the special case of bounded scoring rules, $r_{n}$ will always be of order $\log(n)/{n}^{1/2}$. Further, even for unbounded scoring rules such as the logarithmic scoring rule, the rate $\log(n)/{n}^{1/2}$ is achievable at the expense of additional restrictions and technicalities using techniques such as those in Ghosal and Van Der Vaart (2001). For further discussion and details, see Section 8.
Though Assumption 2 is satisfied by many scoring rules, more interpretable alternatives exist. In particular, Assumption 1 and Jensen’s inequality imply that
                                                                                                    $S\left(P_{Q},x\right)$ ${}={}$ $\int S\left(P_{\theta},x\right)\mathsf{d}Q(\theta)-\Delta(Q,x),$                                                                                                      (7)
for all $x\in\mathcal{X}$ and some non-negative remainder term $\Delta:\mathcal{P}(\Theta)\times\mathcal{X}\to\mathbb{R}_{+}$. In many cases, such as the log predictive score, this remainder does not simplify further. However, kernel scoring rules (Gneiting et al., 2007) can be written as an expectation over a product measure in $\Theta$ for any $x\in\mathcal{X}$ (see Lemma 2 in Appendix B), so that the predictive score is
                                                                                                    $S\left(P_{Q},x\right)$ ${}={}$ $\int\underbrace{\left\{S(P_{\theta_{1}},x_{i})-\delta(x_{i};\theta_{1},\dots,\theta_{k})\right\}}_{\textstyle=:L(\theta_{1:k},x)}\mathsf{d}Q^{k}(\theta_{1:k});\quad Q^{k}(\theta_{1:k})=\bigotimes_{j=1}^{k}Q(\theta_{j}).$                                                                                                      (8)
for some function $\delta:\mathcal{X}\times\Theta^{k}\to\mathbb{R}_{+}$ with $k\in\mathbb{N}_{+}$. This representation allows us to replace Assumption 2 with more interpretable and testable conditions.
Assumption 2.
The predictive scoring rule satisfies (3.1) and for any $\varphi\in\Theta$, if $\theta_{1:k}=(\varphi,\varphi,\dots\varphi)$, then $\delta(x;\varphi,\dots,\varphi)=0$, for all $x\in\mathcal{X}$. Further, for any $\lambda\in\mathbb{R}$, $\mathcal{L}(\theta_{1:k}):=\mathbb{E}[L(\theta_{1:k},X)]$, and some $C>0$, $\mathop{\text{\large$\int$}}\nolimits\mathbb{E}\left[e^{\lambda\{\mathcal{L}(\theta_{1:k})-\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},X_{i})\}}\right]\mathsf{d}\Pi^{k}(\theta_{1:k})\leq e^{\lambda^{2}C^{2}/n}.$
Here, the first part of Assumption 2 is a mild regularity condition, and automatically satisfied by kernel scoring rules for $k=2$. The second part imposes a Hoeffding-type moment condition that generally has to be verified, and which is very similar to those clasically arising in the study of Gibbs posteriors. Though easier to verify, certain scores such as the logarithmic scoring rule violate Assumption 2, so that the more abstract conditions imposed by Assumption 2 must be maintained instead.